# Foundational GPT Model for MEG

### v1

@@ Abstract
Deep learning techniques can be used to first training unsupervised models on large amounts of unlabelled data, before fine-tuning the models on specific tasks. This approach has seen massive success for various kinds of data, e.g. images, language, audio, and holds the promise of improving performance in various downstream tasks (e.g. encoding or decoding brain data). However, there has been limited progress taking this approach for modelling brain signals, such as Magneto-/electroencephalography (M/EEG). Here we propose two classes of deep learning foundational models that can be trained using forecasting of unlabelled MEG. First, we consider a modified Wavenet; and second, we consider a modified Transformer-based (GPT2) model. The modified GPT2 includes a novel application of tokenisation and embedding methods, allowing a model developed initially for the discrete domain of language to be applied to continuous multichannel time series data. We also extend the forecasting framework to include condition labels as inputs, enabling better modelling (encoding) of task data. We compare the performance of these deep learning models with standard linear autoregressive (AR) modelling on MEG data. This shows that GPT2-based models provide better modelling capabilities than Wavenet and linear AR models, by better reproducing the temporal, spatial and spectral characteristics of real data and evoked activity in task data. We show how the GPT2 model scales well to multiple subjects, while adapting its model to each subject through subject embedding. Finally, we show how such a model can be useful in downstream decoding tasks through data simulation. All code is available on GitHub [^2] https://github.com/ricsinaruto/MEG-transfer-decoding.
@@ 1 Introduction
Unsupervised learning provides a dataset-agnostic method for learning shared representations. Unsupervised learning techniques can be further differentiated between those that aim to learn interpretable representations and those that are purely data-driven. In functional neuroimaging, interpretable models can provide neuroscientific insights that are especially useful for rest or spontaneous data, where there is no known external stimuli or behaviour associated with the brain activity (Baker et al.,, 2014). Models designed without focusing on interpretability can be used to generalise over multiple heterogeneous datasets and provide a pretrained foundation model (Yuan et al.,, 2024). Such foundation models can then be fine-tuned for downstream tasks on smaller amounts of data for which relevant labels are available (e.g. to do encoding or decoding). By leveraging large amounts of data, the hope is that the foundation model will be capable of generalising to new data types and provide improvement over just training the model on the labelled dataset. This is especially useful for brain-computer interface (BCI) applications.
The concept of using vast amounts of data to boost performance in downstream tasks originates from deep learning. Perhaps the most successful recent example is that of large language models, trained on diverse data sources and demonstrating enhanced capabilities over task-specific models in a multitude of language-related tasks (e.g., translation, summarisation) (Brown et al.,, 2020). This can also be viewed as a form of transfer learning. Zero-shot performance is obtained when no fine-tuning is done for the downstream task. Several factors enabled the success of large language models, including data scale, model size, fast GPUs, and effective neural network architectures (Kaplan et al.,, 2020; Fedus et al.,, 2022; Sutton,, 2019). To adopt this paradigm for electrophysiology data, the primary obstacles are the model architecture and data size. In this paper, we focus on the former.
We aim to design general models well-suited to multichannel timeseries that can scale effectively. We also focus on foundational models that can be trained using forecasting (as opposed to masked prediction), which are causal and can generate data recursively. This allows for the interrogation of learned spatio-temporal dynamics over long temporal horizons. Self-supervised learning (SSL) has emerged as a promising approach for learning useful representations from unlabelled electrophysiological data. SSL reformulates an unsupervised learning problem as a supervised one by exploiting inherent structure in the data to generate "pseudo-labels". Forecasting is one such SLL task, as is reconstruction of masked sequence segments (Wang et al.,, 2023). In the context of electrophysiology, recent works have proposed SSL tasks tailored to the temporal and multivariate nature of neural time series data (Gohil et al.,, 2022; Banville et al.,, 2021; Kostas et al.,, 2021; Wang et al.,, 2023; Cui et al.,, 2023).
In our quest for designing expressive foundational models of MEG data, we can look to artificial intelligence domains with similar characteristics, such as audio or natural language processing. These domains share some similarities with MEG data, like the sequential nature of the modality. However, while audio data is also a continuous timeseries, it only contains a single channel and comes at a much higher sampling rate compared to M/EEG data. Language data is perhaps even more different as its timeseries are comprised of distinct units (words) from a finite vocabulary set. As such, starting with models developed for these domains and adapting them to handle the nuances of M/EEG data is a promising approach. In this work, we adapt two such approaches. First, we adapt Wavenet, originally developed for forecasting audio data (van den Oord et al.,, 2016); and second, we adapt a Transformer architecture (Vaswani et al.,, 2017) in the form of GPT2, originally developed for forecasting language (Radford et al.,, 2019).
Wavenet has been shown to be an effective model for forecasting time series, through its dilated convolutional architecture, which is fast and parameter-efficient (van den Oord et al.,, 2016). Here, we modify Wavenet to handle multichannel MEG data by first doing a channel-dependent quantisation (WavenetFullChannel) and then introducing a mixing layer across the channel dimension (WavenetFullChannelMix).
In recent years, the Transformer architecture has driven a second deep learning revolution (Vaswani et al.,, 2017). Transformers use attention to model complex dependencies in long sequences, providing a more flexible inductive bias well-suited to language modelling and other tasks involving highly structured sequential data (Devlin et al.,, 2019; Brown et al.,, 2020). For example, self-attention provides direct connectivity between any two time steps, capturing long-range dependencies. The parallelisable architecture allows for more efficient computation compared to recurrent models. Transformers have also been applied to time series data with promising results (Wen et al.,, 2022). Zhou et al., (2021) adapted the self-attention mechanism for long-range forecasting and demonstrated state-of-the-art performance on multiple public datasets.
Here, we modify the Transformer architecture used in GPT2 (Radford et al.,, 2019) to handle multichannel MEG data. Specifically, we do a channel-independent quantisation, but augment the input with channel-embeddings to signify to the model which timeseries corresponds to which channel. We refer to this model as ChannelGPT2. While previous work has employed various modifications to the Transformer architecture we posit that keeping the categorical nature of the sequences and doing next time-step forecasting are essential for effective modelling. Our methods are further detailed in Section 3.
Both the modified Wavenet and ChannelGPT2 models also include subject embeddings (Csaky et al., 2023b, ) and task condition embeddings. The latter provides the model with information about external stimuli synchronised to the MEG time-course. This means that the model can be treated as both a forecasting and an encoding model.
We evaluate how well the modified Wavenet and ChannelGPT2 models can perform as foundational models. First, we assess how well the fitted models can recursively generate data with the same spatial, temporal and spectral characteristics as real MEG data. We evaluate this both at the subject- and group-level. Second, we see how well the foundation models can be used on downstream tasks, through data simulation. Finally, we show through several ablation studies how the best performing model, ChannelGPT2, leverages condition and channel embeddings.
@@ 2 Results
2.1 Next time-step prediction does not capture performance
@@
We wanted to evaluate how well the modified Wavenet and ChannelGPT2 models can perform as foundational models. One way to assess this is through the forecasting performance, i.e. the prediction accuracy of the label at the next time point. Here, we evaluated the performance of our two modified versions of Wavenet (WavenetFullChannel and WavenetFullChannelMix) alongside ChannelGPT2, as summarised in Table 1. For comparison, we also evaluated the performance of a linear autoregressive (AR) model of order 255. All deep learning models were trained on tokenised and embedded inputs, and cross-entropy loss was used to predict categorical probability distributions over tokens. AR models were trained on the continuous data using the mean-squared error loss, and we simply binned the predicted continuous output to compute accuracy and compare with other models.
@@
The next-timestep forecasting accuracy for different models on a sample subject is shown in Tabel 2. Beyond standard accuracy (the number of true positives divided by the number of all examples), we also evaluated top-5 accuracy, counting a prediction as correct if the true bin was within the 5 most probable bins. Surprisingly, all models performed only slightly better than a naive baseline of repeating the previous timestep’s value.
As expected, the linear AR model had lower MSE but worse accuracy than the nonlinear models. This can be because MSE measures the distance of the prediction to the target, while accuracy is only 1 if the prediction is in the target bin. Thus, it may be that the AR model always predicts values that are slightly closer to the target, but never quite falling in the target bin. While WavenetFullChannel appears to be worse, WavenetFullChannelMix and ChannelGPT2 have nearly identical performance.
The choice of sampling rate can affect forecasting performance. A higher sampling rate makes the task easier as consecutive timesteps are more correlated, however this might make the model focus on very short-range temporal dependencies and overfit to noise. We analysed forecasting performance in relation to sampling rate in Supplementary Section A.2.
Model Univariate Tokenised Linear
AR(255) yes no yes
WavenetFullChannel yes yes no
WavenetFullChannelMix no yes no
ChannelGPT2 yes yes no
Table 1: The 4 main models presented in the results section. Univariate means that the channel dimension is treated as a batch dimension during training. Non-tokenised models predict continuous values directly instead of categorical distributions.
Model MSE Top-1 Accuracy Top-5 Accuracy
Repeat baseline 0.024 1.5 7.6
AR(255) 0.016 1.5 7.5
WavenetFullChannel 0.026 2.0 9.8
WavenetFullChannelMix 0.022 2.2 10.8
ChannelGPT2 0.023 2.2 10.9
@@
Table 2: Test data next-timestep prediction performance across various models. Accuracy values are given in percentages. Chance-level is $1/256$, however predicting the majority class (quantised value) is somewhat higher, since the distribution over bins is not uniform.
2.2 PSD of generated data matches real data
We have seen how forecasting accuracy does not differentiate well between the candidate foundation models. Despite this, the models may perform differently when evaluated using other criteria. For example, a good foundational model should be expected to be able to recursively generate data that looks like the real data. Here, we first assess the models’ ability to do this using the power spectra.
@@ 2.2 PSD of generated data matches real data
For deep learning models we used top-p sampling with $p=80\%$ (unless otherwise noted in the figure caption) to recursively generate data. We generated 3600 seconds with all models. For models that have task-conditioning (all except AR(255)) we use the task label timeseries from the training set. The models in this section were trained on a single sample subject.
@@
(c) WavenetFullChannel
(d) ChannelGPT2
(e) WavenetFullChannelMix p=0.72
(f) WavenetFullChannelMix p=0.80
Figure 1: Comparison of generated data power spectral density (PSD) across models. Each line represents a different MEG channel. Note that in e) and f) the value of p corresponds to the probability mass used for top-p sampling.
@@
Generated token sequences are first de-tokenised and then the power spectral density (PSD) is computed on the continuous data. Figure 1 compares the PSD of the generated data across our models. Qualitatively, it is clear that AR(255) reproduces PSDs that match best with those computed directly on the MEG data, while WavenetFullChannel and ChannelGPT2 are not far behind. All models capture the characteristic $1/f$ shape, and peaks at 10 and 19 Hz, likely related to alpha and beta band activity. Notably, WavenetFullChannel has reduced power at the 19 Hz peak, which could indicate issues in capturing higher frequency dynamics.
The performance of WavenetFullChannelMix was found to depend on the choice of the probability mass $p$, used in top-p sampling from the model to generate the data. $p$ prohibits the generation of any tokens that are not within the top-p cumulative probability mass. Even slight modifications (e.g., $p=0.72$ vs. $p=0.8$ for WavenetFullChannelMix) result in large differences in the frequency of the two peaks in the PSDs computed on the generated data, and also the width of the peaks. Ultimately both top-p values provide subpar PSDs compared to channel-independent models. This is likely to be due to overfitting, as channel-mixing models lack the implicit regularisation of modelling each channel separately. The implicit regularisation is a consequence of having much more examples and much fewer input features when treating channels as a batch dimension.
2.3 HMM statistics of generated data match real data

### v2

# GPT2MEG: Quantizing MEG for Autoregressive Generation

@@ Abstract
Foundation models trained with self-supervised objectives are increasingly applied to brain recordings, but autoregressive generation of realistic multichannel neural time series remains comparatively underexplored, particularly for Magnetoencephalography (MEG). We study (i) modified multichannel WaveNet variants and (ii) a GPT-2-style Transformer, autoregressively trained by next-step prediction on unlabelled MEG. For the Transformer, we propose a simple quantization/tokenization and embedding scheme (channel, subject, and task-condition embeddings) that repurposes a language-model architecture for continuous, high-rate multichannel time series and enables conditional simulation of task-evoked activity. Across forecasting, long-horizon generation, and downstream decoding, GPT2MEG more faithfully reproduces temporal, spectral, and task-evoked statistics of real MEG than WaveNet variants and linear autoregressive baselines, and scales to multiple subjects via subject embeddings. Code available at https://github.com/ricsinaruto/MEG-transfer-decoding.
Magnetoencephalography, Time series, Autoregressive modeling, Generative modeling, GPT, Tokenization, Conditional generation, multi-subject modeling
@@ 1 Introduction
Magnetoencephalography (MEG) and related electrophysiology (EEG/ECoG) provide millisecond-resolved measurements of large-scale brain dynamics, but remain challenging to model due to high dimensionality, strong temporal structure, and substantial variability across subjects and recording setups. Recent self-supervised foundation models have begun to learn transferable representations from large collections of unlabelled neural recordings, enabling improved performance and data efficiency on downstream decoding/encoding tasks (Kostas et al., 2021; Wang et al., 2023; Cui et al., 2023; Jiang et al., 2024; Wang et al., 2024; El Ouahidi et al., 2025). Most existing brain foundation models focus on representation learning via masked or contrastive objectives, producing embeddings for classification and regression. In contrast, autoregressive generative foundation models, capable of simulating realistic neural time series and supporting data augmentation, uncertainty-aware decoding, and mechanistic hypothesis testing, have received comparatively less attention, especially for MEG.
Autoregressive sequence models are a natural fit for neural time series. Going back to the pre-foundation-model era, in audio, WaveNet modeled waveforms via causal dilated convolutions over discretized samples (van den Oord et al., 2016). In language, decoder-only Transformers learn long-range dependencies through self-attention over discrete tokens (Vaswani et al., 2017; Radford et al., 2019). Inspired by recent work on tokenized time-series models (Ansari et al., 2024; Das et al., 2024; Rasul et al., 2023), we ask: can we repurpose autoregressive language-model training to build a generative ”foundation” [^1] While our experiments are small-scale, the framework has the potential to scale to produce foundational models. model for multichannel MEG?
We propose two classes of autoregressive models trained purely by next-step prediction on unlabelled MEG: (i) modified multichannel WaveNet variants and (ii) a modified GPT-2 model we call GPT2MEG. Our key contribution is a simple quantization/tokenization and embedding scheme that allows a discrete-token Transformer to model continuous multichannel MEG. We further incorporate time-aligned task-condition labels as local inputs, enabling conditional generation of task-evoked responses, and we scale to multiple subjects via subject embeddings (Csaky et al., 2023a).
Contributions.
  • We introduce GPT2MEG, an autoregressive Transformer model for sensor-level MEG using discrete tokenization with channel-, subject-, and stimulus-embeddings for conditional generation.
  • We benchmark Transformer- and WaveNet-based generative models against a linear AR baseline, showing that one-step forecasting accuracy is insufficient; instead we evaluate long-horizon simulations via spectral metrics, HMM-derived dynamics, and task-evoked responses.
  • We demonstrate multi-subject scaling with subject embeddings and show that simulated trials from a multi-subject model can improve downstream decoding through transfer learning.
2 Related Work
Self-supervised and foundation models for electrophysiology. Self-supervised learning has enabled representation learning from unlabelled EEG/MEG using contrastive, masked, and predictive objectives (Banville et al., 2021; Kostas et al., 2021; Wang et al., 2023). Recent work scales these ideas towards foundation models for brain signals (Cui et al., 2023; Jiang et al., 2024; Wang et al., 2024; El Ouahidi et al., 2025). Transformers are increasingly used in EEG analysis; for a recent review see Vafaei and Hosseini (2025). For MEG, related sequence models have also been explored for stimulus-to-brain encoding (Chehab et al., 2022). These approaches primarily target embedding learning for decoding/encoding; our focus is complementary: building autoregressive generative models that can simulate realistic multichannel MEG and task-evoked activity.
Generative models for brain signals. Generative modelling of electrophysiology has been explored with VAEs, GANs, diffusion models, and Transformers, often motivated by data augmentation in BCI settings. Recent examples include transformer-based EEG synthesis with learned discrete codes (Lim and Kuo, 2024). We study sensor-level MEG with explicit channel, subject, and task-condition embeddings, and we evaluate generation using multivariate dynamical metrics (HMM state statistics) in addition to spectral and evoked-response analyses.
Tokenized time-series language models. Several recent foundation models for generic time-series forecasting tokenize continuous values into discrete vocabularies and train language-model architectures with cross-entropy loss, enabling zero-shot or few-shot generalization (Ansari et al., 2024; Das et al., 2024; Rasul et al., 2023). Our GPT2MEG adopts a similarly simple discretization ($\mu$-law companding + uniform bins) but addresses challenges specific to MEG: high-dimensional sensor arrays, subject variability, and time-aligned task conditioning.
3 Methods
3.1 Problem setup and notation
We consider multichannel sensor-level MEG recordings $\mathbf{X}\in\mathbb{R}^{C\times T}$ with $C$ sensors and $T$ time points. After preprocessing and scaling, each sample is discretized into one of $Q$ tokens via $\mu$-law companding followed by uniform quantization (Section 3.2). For task datasets we also have a time-aligned condition sequence $\mathbf{y}_{1:T}$ (e.g. stimulus identity), and for multi-subject training a subject index $s$. Our goal is to learn a causal conditional generative model
$p(\mathbf{X}\mid\mathbf{y},s)=\prod_{t=1}^{T}p(\mathbf{x}_{t}\mid\mathbf{x}_{1:t-1},\mathbf{y}_{\leq t},s),$ (1)
where $\mathbf{x}_{t}\in\mathbb{R}^{C}$ is the multichannel sample at time $t$. Some of the models we evaluate factorize across sensors (treating the channel dimension as a batch dimension during training), while others include explicit mixing across channels. All deep learning models are trained by teacher-forced next-token prediction with cross-entropy loss. We evaluate both one-step forecasting and long-horizon recursive generation using spectral, multivariate-dynamical, and task-evoked metrics (Section 3.4).
3.2 Multi-channel Wavenet
We adapt WaveNet (van den Oord et al., 2016) as an autoregressive baseline for multichannel MEG. The amplitude of each channel (and timestep) is discretized independently into Q=256Q{=}256 bins using $\mu$-law companding followed by uniform quantization. Concretely, after scaling each channel to (1,1)(-1,1) we apply
$f(x)=\mathrm{sign}(x)\,\frac{\ln(1+\mu|x|)}{\ln(1+\mu)},\quad\mu{=}255,$ (2)
and then uniformly bin $f(x)$ into $Q$ discrete tokens. Training with cross-entropy enables sampling from a learned categorical distribution and helps avoid mean-prediction bias from MSE regression (Banville et al., 2021).
We evaluate two variants. WavenetFullChannel (WFC) treats channels as a batch dimension: one shared causal dilated-convolution stack is applied independently to each channel, with channel identity provided via a per-channel embedding of the discrete tokens. WavenetFullChannelMix (WFCM) additionally applies a learned $C\times C$ mixing projection to the skip representation to share information across channels. For task data we optionally provide time-aligned condition labels (and subject IDs for group training) through standard WaveNet local conditioning embeddings. Full equations are in Appendix A.1.
3.3 Multi-channel GPT2
We set out to design a Transformer model suited for M/EEG data, while keeping the key elements that made it successful in language modelling. Specifically, we use GPT-2. When adapting it to continuous multivariate time series, the main challenges are at the input and output layers interfacing the model with the data.
To apply GPT2 to our continuous multichannel time series data, we take a similar approach as with Wavenet by tokenising each channel independently using the same method as before. This serves as our equivalent of the discrete set of tokens in language modelling. The same GPT2 model is applied to each channel in parallel by setting the channel dimension as the batch dimension. We call this GPT2MEG.
The input to the model includes the position embedding as well as subject and task-stimulus embeddings. We also add a label/embedding telling GPT2 which channel the current time series is coming from:
where ++ denotes element-wise addition, $\mathbf{X}\in\mathbb{R}^{C\times T\times Q}$ is the tokenised input, $\mathbf{W}_{c}\in\mathbb{R}^{C\times E}$ are the learned channel embeddings of size $E$, which are distinct for each channel $c\in{1,\dots,C}$ but constant across time $t$. $\mathbf{Y}$ and $\mathbf{O}$ are the task and subject index matrices, mapped to their respective embeddings with $\mathbf{W}_{y}\in\mathbb{R}^{Y\times E}$ and $\mathbf{W}_{o}\in\mathbb{R}^{O\times E}$. $Y$ is the number of conditions and $O$ is the number of subjects. During timesteps when there is no stimulus present, $\mathbf{Y}\mathbf{W}_{y}$ is set to zero. As with the positional encoding $\mathbf{W}_{p}$, we simply add all embeddings (task, subject, channel) into a single representation. Note that instead of having channel-specific embeddings of the tokenised input $\mathbf{X}$ we learn the same mapping $\mathbf{W}_{e}\in\mathbb{R}^{Q\times E}$ across channels. Channel information is provided to the model through the channel embeddings.
3.4 Generation and evaluation
We evaluate models in both forecasting and free-running generation settings. Tokenized models (WaveNet and GPT2) generate by sampling tokens autoregressively (top-$p$ sampling (Holtzman et al., 2020)); AR models generate by recursively filtering noise. Since one-step prediction accuracy can be weakly informative for long-horizon behaviour, we emphasize generation-based metrics:
  • Spectral fidelity: power spectral density (PSD) comparisons.
  • Multivariate dynamics: summary statistics of a 12-state time-domain embedding HMM fit to generated vs real multichannel time series (Vidaurre et al., 2018b).
  • Task-evoked structure: trial-averaged evoked responses under task conditioning.
  • Downstream utility: decoding performance when training/fine-tuning classifiers on simulated trials.
Implementation details for HMM fitting and decoding are provided in Appendix A.2.
4 Experimental Setup
Dataset. We use the 15-subject continuous MEG dataset of Cichy et al. (2016) (118 visual images, 30 trials/image). We apply standard preprocessing (filtering, ICA artifact rejection) and downsample to 100 Hz; details are in Appendix A.3. For within-subject trainings we create non-overlapping train/validation/test splits by holding out 4 trials/condition for validation and 4 for test (remaining 22 for training).
Tokenization. Each sensor is standardized (0 mean, unit variance), clipped (10,10)(-10,10), rescaled to (1,1)(-1,1), and quantized into Q=256Q{=}256 bins with $\mu$-law companding. This discretization incurs negligible loss for evoked and decoding analyses (Appendix A.4).
Models and training. We compare a per-channel linear AR(255) baseline (order matched to the deep models’ receptive field) with WFC, WFCM, and GPT2MEG. Unless otherwise stated, generative analyses are reported for a representative subject (single-subject training) to enable rapid iteration, and we additionally evaluate multi-subject scaling with GPT2MEG-group (15 subjects). WFC uses a receptive field of 255 samples via two stacks of dilated-convolution blocks; GPT2MEG uses 12 decoder layers, 12 heads, and $d_{\mathrm{model}}{=}96$ with variable context length (128–256). Deep models are trained with cross-entropy on quantized tokens and early stopping on validation loss; AR is trained with MSE on continuous signals. Full hyperparameters are reported in Appendix A.5.
@@ 5 Results
5.1 One-step prediction is weakly informative
@@
First, we assessed the models’ forecasting performance, i.e. the prediction accuracy of the label at the next time point. We used two different modified versions of Wavenet (WavenetFullChannel and WavenetFullChannelMix) alongside GPT2MEG. For comparison, we also evaluated the performance of a linear autoregressive (AR) model of order 255. For AR models we simply binned the predicted continuous output to compute accuracy and compare with other models.
@@
The results on a sample subject is shown in Figure 1. Beyond standard accuracy, we also evaluated top-5 accuracy, counting a prediction as correct if the true bin was within the 5 most probable bins. Surprisingly, all models performed only moderately better than a naive baseline of repeating the previous timestep’s value. However, as we shall investigate later, this does not necessarily reflect the richness of the structure in data recursively generated by the models.
@@
Figure 1: Next-timestep prediction performance across the different forecasting models. Accuracy values are on left-out test data and are given in 0-1 units. Chance-level is $1/256$, however predicting the majority class (quantised value) is somewhat higher, since the distribution over bins is not uniform. ChannelGPT refers to GPT2MEG.
As expected, the linear AR model had lower MSE but worse accuracy than the nonlinear models. This can be because MSE measures the distance of the prediction to the target, while accuracy is only 1 if the prediction is in the target bin. Thus, it may be that the AR model always predicts values that are slightly closer to the target, but never quite falling in the target bin. While WavenetFullChannel appears to be worse, WavenetFullChannelMix and GPT2MEG have nearly identical performance. Based on these results it is inconclusive whether deep learning models improve over the linear AR model. Further, long-range structure analyses are presented in the next sections to elucidate this.
We note that the choice of sampling rate can affect forecasting performance. A higher sampling rate makes the task easier as consecutive timesteps are more correlated, however this might make the model focus on very short-range temporal dependencies and overfit to noise. We analysed forecasting performance in relation to sampling rate in Supplementary Section A.9.
5.2 Spectral fidelity of long-horizon generation
@@
A good generative model should be expected to be able to recursively generate data that looks like the real data. Here, we first assess the models’ ability to do this using the power spectra. For deep learning models we used top-p sampling with $p=80\%$ (unless otherwise noted in the figure caption) to recursively generate data. We generated 3600 seconds with all models. For models that have task-conditioning (all except AR(255)) we use the task label timeseries from the training set. The models in this section were trained on a single sample subject, containing about 1.5 hours of data downsampled to 100 Hz across 306 channels.
@@
(c) WFC
(d) GPT2MEG
Figure 2: PSD comparison between real MEG (a) and long-horizon generations from different models (b–d) for a representative subject. Each line is a sensor/channel. Additional WFCM sampling results, generally producing less realistic spectra, are in Appendix A.5.
@@
Generated token sequences are first de-tokenised and then the power spectral density (PSD) is computed on the continuous data. Figure 2 compares the PSD of the generated data across our models. Qualitatively, it is clear that AR(255) reproduces PSDs that match best with those computed directly on the MEG data, while WavenetFullChannel and GPT2MEG are not far behind. However, a more detailed analysis on HMM power spectra (Figure 9 in the Appendix) shows that the heterogeinity of frequency content is not capture by the AR model. Thus, while average PSD across time may look great, it cannot be used as a single evaluation point. All models capture the characteristic $1/f$ shape, and peaks at 10 and 19 Hz, likely related to alpha and beta band activity. Notably, WavenetFullChannel has reduced power at the 19 Hz peak, which could indicate issues in capturing higher frequency dynamics. Generated PSDs of group-level models trained on all subjects (Section 5.5) were equally similar to the real data (plots not shown).
5.3 Transformer generation better matches multivariate dynamics